600.0 mL of H2SO4 is titrated with a 6.0 M solution of NaOH and requires only 6.8 mL of NaOH to reach the endpoint. According to the unbalanced equation H2SO4 + NaOH → H2O + Na2SO4, what is the concentration of the H2SO4?

A. 0.041 moles
B. 0.021 moles
C. 0.35 M
D. 0.035 M
E. 0.0035 M
1
Expert's answer
2018-03-08T10:59:07-0500
We write the equation of the reaction, setting the coefficients:
H2SO4 + 2NaOH = Na2SO4 + 2H2O
By the condition of the problem, 6.8 ml of 6M sodium hydroxide solution was required for acid titration, then the number of moles of sodium hydroxide reacted would be:
n(NaOH)=c(NaOH)*V(solution)=6*0.0068=0.0408 moles
By the reaction equation: n(H2SO4)=1/2×n(NaOH)=0.0204 moles
Since the volume of the solution of titrated acid is known equal to 600 ml (0.6 L), we find its concentration: с(Н2SO4)=(n(H2SO4))/(V(solution))=0.0204/0.6=0.035 M